OpenStudy (anonymous):

sin[sin3+sin5]=cos[cos3-cos5] Establish the identity I used the sum to product formula so far

OpenStudy (jdoe0001):

is that a literal 3 and 5? or are those exponents?

OpenStudy (jdoe0001):

that is \(\bf \large sin[sin(3)+sin(5)]=cos[cos(3)-cos(5)] ?\)

OpenStudy (anonymous):

yes it's just three and five, they have thetas after them as well @jdoe0001

OpenStudy (jdoe0001):

hmmm

OpenStudy (anonymous):

I have to show how the left side equals the right without changing the right hand side

OpenStudy (anonymous):

Using trig functions

OpenStudy (jdoe0001):

well... if you notice the graph you'd end up in dead ends though

OpenStudy (asnaseer):

Is this what you are trying to show?\[\sin(\theta)\left(\sin(3\theta)+\sin(5\theta)\right)=\cos(\theta)\left(\cos(3\theta)-\cos(5\theta)\right)\]

OpenStudy (anonymous):

Yes

OpenStudy (asnaseer):

OK - notice how in this identity we have \(\theta,3\theta,5\theta\) where we notice that:\[\theta+3\theta=4\theta\]and:\[5\theta-\theta=4\theta\]

OpenStudy (asnaseer):

Can you see how to make use of this fact?

OpenStudy (asnaseer):

Think about the cosine addition and subtraction formulas

OpenStudy (anonymous):

I tried using the sum to product formula

OpenStudy (asnaseer):

From what I showed you above we can write:\[\theta+3\theta=5\theta-\theta\]Now just take cosines of both sides to get:\[\cos(\theta+3\theta)=\cos(5\theta-\theta)\] Finally expand each of these using the cosine addition/subtraction formulas, rearrange and you will get to the desired identity.

OpenStudy (anonymous):

I dont understand how you got that in the previous part @asnaseer

OpenStudy (asnaseer):

1+3=5-1=4

OpenStudy (asnaseer):

Do you understand?

OpenStudy (anonymous):

Isnt the sin on the outside multiplied

OpenStudy (asnaseer):

Yes it is, and if you expand this out as I suggested then you will get to your identity:\[\cos(\theta+3\theta)=\cos(5\theta-\theta)\]

OpenStudy (asnaseer):

Have you been asked to specifically use the sum to product formulas?